# CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation

### v2

@@ Abstract
Machine Learning, ICML
1Beihang University, 2Tsinghua University, 3National University of Singapore
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation. While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge (Liu et al., 2024; Zhang et al., 2023). To drive progress in this field, a comprehensive and systematic evaluation framework is essential (Jiang et al., 2024; Zhou et al., 2023).
This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation. Built upon insights from existing evaluation sets like McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++. It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
To ensure rigorous assessment, we propose four novel evaluation metrics: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics measure models’ ability to handle multi-constraint problems from different perspectives, including full compliance, average constraint satisfaction, logical coherence, and consistency in instruction execution. By offering a structured evaluation framework, CodeIF provides valuable insights into the current state and future direction of LLM-driven code generation.
Overall, our contributions are mainly four-fold:
@@
    Innovative Benchmark: We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation. CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
@@
    Automated High-Quality Instruction Generation: Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists. This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
@@
    Novel Evaluation Metrics: We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks. These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
@@
    Extensive Evaluation and Analysis: We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels. Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
@@ 2 CodeIF
As shown in Figure 1, the construction of CodeIF includes the following steps: we first collect and refine constraint instructions from real code generation tasks, then we combine code generation tasks with LLMs and human review to create a high-quality evaluation dataset.
@@ 2.1 Building
The construction of the CodeIF dataset consists of two main phases: collection of constraint instructions and data processing. In Section  2.2, we elaborate on how we collected the constraint instructions for code generation. Following this, in Section  2.3, we describe how we combined these constraint instructions to produce the final CodeIF evaluation dataset.
@@ 2.2 Constraint Instructions Collection
The first phase of our work focuses on code generation by collecting and formulating constraint instructions to construct the CodeIF evaluation dataset. This process consists of two stages: (1) collecting and verifying constraint instructions and (2) applying them to dataset generation.
We begin with an in-depth analysis of code generation tasks based on existing benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), establishing an instruction system covering eight categories. Each category targets a specific aspect of code generation, enabling a fine-grained assessment of LLMs’ instruction-following abilities.
Within each category, we further refine constraints into atomic instructions with clear, explicit directives. These atomic constraints allow binary evaluation (“yes” or “no”), ensuring an objective and precise assessment of model compliance while minimizing subjectivity.
Our classification of atomic constraint instructions is guided by an analysis of existing code generation benchmarks and widely used code repositories. We observe that requirement distribution varies significantly across different levels—line-level, function-level, and full-solution generation. To ensure comprehensive coverage, we design atomic instruction categories that assess both high-level architectural constraints and fine-grained variable control.
To systematically evaluate LLMs’ instruction-following capabilities in code generation, we define eight constraint categories. The Global category ensures that generated code aligns with overarching specifications and holistic constraints. Structural Control focuses on adherence to control structures such as loops and conditionals, as well as data structures like sets and maps. Variable constraints assess details at the variable level, including naming conventions and initialization methods.
At different abstraction levels, Interface, Function, and Class constraints evaluate the model’s ability to construct well-defined program components. The File category tests handling of cross-file dependencies and external library calls. Finally, the Combination category presents a comprehensive challenge by integrating constraints from multiple dimensions, assessing how well models follow complex, multi-faceted instructions in realistic scenarios.
@@
As shown in Figure 2, CodeIF’s distribution across different programming languages and categories is presented.
Through a detailed analysis of existing code generation benchmarks, we designed an evaluation system comprising 8 main categories and 50 fine-grained constraint instructions. This system not only enables a systematic assessment of LLMs’ performance in code generation but also establishes clear benchmarks for future research.
These instructions cover constraints ranging from high-level structural guidelines to fine-grained variable operations, ensuring a comprehensive evaluation. For instance, the Global category assesses a model’s ability to maintain overall structural integrity and logical coherence, while the Variable category evaluates its precision in variable naming and initialization.
By refining constraints into well-defined categories, we can precisely identify the strengths and weaknesses of different LLMs, guiding optimization efforts and advancing the field toward more effective automated code generation. Additionally, this structured evaluation approach provides a valuable reference for future research, fostering continuous innovation. The complete list of atomic constraint instructions is provided in Appendix Table 4.
@@ 2.3 Data Construction
To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024). These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
@@
To construct high-quality constraint instructions, we leveraged large language models (LLMs) such as GPT-4 to automatically generate task-specific instruction lists. Initially, we curated approximately 20 detailed examples, each featuring a specific code generation task. Human experts then formulated atomic constraint instructions to ensure precise adherence during code generation.
These examples served as prompts, allowing LLMs to generate corresponding instruction lists while refining task descriptions by identifying and eliminating redundant constraints. This approach streamlined instruction-following requirements, ensuring clarity and promoting high-quality outputs.
@@
Since some atomic constraints exhibit interdependencies, we further utilized LLMs to construct explicit dependency relationships within instruction sets. These dependencies were integrated into our evaluation framework, enhancing assessment depth and improving the accuracy and practicality of constraint verification.
Specifically, by analyzing the logical connections and sequential dependencies among atomic instructions, we can precisely define the execution steps required for each task. For instance, in function generation tasks, the process typically involves defining the function name first, specifying parameter types, and finally implementing the function body. These dependencies ensure coherence and consistency throughout the code generation process.
Integrating these dependencies into the evaluation system enables a more accurate assessment of a model’s ability to follow complex instructions while providing a foundation for future optimizations. Additionally, this approach helps identify potential challenges models face when handling specific instruction sequences, offering insights for further improvement. As shown in Figure 3, this example from CodeIF illustrates a specific generation task, its associated instruction list, and the corresponding instruction dependencies.
@@ 2.4 Data Analysis
As shown in Table 1, the dataset is categorized into three difficulty levels: Easy, Hard, and an aggregated set, Full. The Easy and Hard sets each contain 600 code generation tasks, while the Full dataset combines both, totaling 1,200 tasks. These tasks span four programming languages: Go, Python, Java, and C++.
Among them, Java has the highest number of tasks (353), followed by Python (348). C++ and Go contain 269 and 230 tasks, respectively. The dataset also records the average number of instructions per task for each difficulty level. The Easy set averages 11.99 instructions per task, while the Hard set increases to 13.8, reflecting greater complexity and stricter constraints. The Full dataset, combining both, has an average instruction length of 12.9. Figure 4 illustrates the distribution of task lengths across the dataset.
@@
Analysis of Constraint Instruction Category Distribution
Table 2 presents the distribution of atomic instructions across different difficulty levels in the evaluation set. As shown in Figure 5, we compare the average number of instructions per category between the “Easy” and “Hard” sets. These categories include Global, Structural Control, Variable, Interface, Function, Class, File, and Combination.
Across all categories, the “Hard” dataset contains a higher average number of instructions than the “Easy” set. Notably, the Global category has the highest instruction count in both difficulty levels, averaging 2.5 instructions in the “Easy” set and exceeding 3 in the “Hard” set. This trend indicates that as task complexity increases, models face greater challenges in handling intricate constraints and detailed requirements. By analyzing instruction distribution across categories, we gain deeper insights into model performance under varying levels of difficulty, providing a foundation for further optimization.
Figure 6 illustrates the proportion of each atomic instruction category. The Global category holds the largest share at 22.77%, followed by the Variable category at 18.17%. This distribution highlights CodeIF’s focus on evaluating LLMs’ capabilities at both the overall structural level (e.g., Global constraints) and fine-grained variable handling (e.g., Variable constraints). This design ensures a balanced assessment of code generation, emphasizing both high-level logical coherence and precision in variable management.
@@ 3 Metrics
Completely Satisfaction Rate (CSR)
@@
where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied. CSR measures the proportion of problems where all constraints are fully met.
Soft Satisfaction Rate (SSR)
@@
where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on. RSR ensures that models satisfy prerequisite constraints before fulfilling dependent ones.
@@
$\small\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$. CCSR evaluates a model’s consistency in following sequential instructions without errors.
@@ 4 Experiment
4.1 Baselines
We evaluate over 30 language models, including both open-source models and commercial API services.
  • Meta Llama 3 Series (Touvron et al., 2023): Includes Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct.
  • Qwen2.5 Series (Yang et al., 2024): Includes Qwen2.5-1.5B-Instruct, Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct. Additionally, it includes dedicated code generation models: Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024), Qwen2.5-Coder-3B-Instruct, Qwen2.5-Coder-7B-Instruct, Qwen2.5-Coder-14B-Instruct, and Qwen2.5-Coder-32B-Instruct.
  • Mistral Series (Jiang et al., 2023): Includes Mistral-3B, Mistral-8B, and the specialized code model Codestral-2501.
  • Microsoft Phi Series (Abdin et al., 2024): Includes Phi-3.5-Mini-128K-Instruct (3.8B) and the latest Phi-4.
  • Google Gemma Series (Team, 2024b): Includes Gemma-2-9B-It and Gemma-2-27B-It.
  • DeepSeek Series: Includes DeepSeek-Coder (Guo et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024).
  • OpenAI Series (Achiam et al., 2023): Includes GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and the latest GPT-4O-2024-11-20.
  • Google Gemini Series (Team, 2024a): Includes Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro.
  • Anthropic Series: Includes Claude-3.5-Sonnet-20241022.
@@ 4.2 Experimental Setup
The temperature coefficient is set to 0 to ensure output determinism, with a maximum generation length of 4096 tokens. All other settings follow the official default parameters for each model. Commercial API models are accessed through the latest available interface as of December 2024. Specifically, this study evaluates the Qwen2.5-Coder series, Meta’s latest Llama 3.3 architecture, and the DeepSeek-V3 large-scale dialogue model as upper-bound benchmarks. All experiments are conducted using the official API and 8 H800(80G).
@@ 4.3 Automatic Evaluation
To ensure robust evaluation, we employed both large language models (LLMs) and human experts to assess whether each model correctly followed the atomic constraint instructions. During data preprocessing, all constraints were decomposed into atomic elements, allowing the adjudication system to perform objective binary evaluations (Yes/No) instead of relying on subjective judgment.
Follow by FairEval (Wang et al., 2023a), GPT-4-1106-Preview model served as the primary automated evaluation tool (see Appendix 8.1 for the detailed prompt). Additionally, three domain experts manually annotated a stratified sample of 150 instances. Statistical analysis revealed strong agreement between LLM-based and human evaluations, with Pearson correlation coefficients of 0.87 (LLM-human) and 0.85 (inter-human), demonstrating high consistency between automated and manual assessments.
These results suggest that the atomic constraint decomposition methodology effectively reduces subjectivity, making LLM-based evaluation a reliable and scalable approach for constraint verification tasks.
@@ 4.4 Main Experiments
Table 3 provides a comprehensive evaluation of CodeIF. The results are measured using four key metrics: Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). More detailed results are in Appendix 8.2.
Result Overview
DeepSeek-V3 and Claude-3-5-Sonnet-20241022 consistently achieve top scores across all difficulty levels and evaluation metrics, excelling in complex tasks. However, even the highest CSR on the Hard dataset is only 0.362, highlighting room for improvement in handling strict constraints.
Model Scale and Performance Trends
Larger models generally perform better across all metrics. In the Qwen2.5 series (Qwen2.5-1.5B-Instruct to Qwen2.5-72B-Instruct), performance improves with model size. However, this trend is not universal. In the Llama3 series, larger models do not always achieve the best results, suggesting that architecture design, training data quality, and optimization strategies are equally critical factors.
Open-Source vs. Closed-Source Models
Closed-source models like GPT-4O and Claude-3-5-Sonnet-20241022 outperform their open-source counterparts, maintaining higher accuracy under complex constraints. While some large open-source models (e.g., Qwen2.5-72B-Instruct) show competitive results, they still lag behind. This performance gap likely stems from advantages in data quality, optimization techniques, and reinforcement learning from human feedback (RLHF), which proprietary models leverage for enhanced instruction-following precision.
Impact of Task Difficulty
Table 3 shows that as task complexity increases (from Easy to Hard), the performance of nearly all models declines. For example, GPT-4O-2024-11-20 achieves a CSR of 0.441 on Easy tasks, but this drops to 0.325 on Hard tasks. This trend confirms that higher constraint levels impose greater challenges, reducing instruction-following accuracy. These findings align with previous research, which suggests that models’ ability to adhere to instructions deteriorates as task difficulty increases.
@@ 5.1 Performance Analysis Across Instruction Types
Figure 7 illustrates the performance of LLMs across different instruction categories, revealing notable disparities among models. DeepSeek-V3 consistently achieves the highest scores across most evaluation dimensions, with a particularly strong performance in combination tasks (0.831), highlighting its effectiveness in handling complex instructions. However, its strengths are not uniform; it excels in global structure control but shows relative weaknesses in variable handling, likely reflecting its specific optimization strategies.
Conversely, Meta’s Llama series demonstrates a significant correlation between model size and performance. While smaller configurations (Llama-3.2-1B-Instruct) score lower, larger variants (Llama-3.3-70B-Instruct) perform significantly better, showcasing substantial gains with increased parameter counts. However, model size alone is not the sole determinant of performance. Comparisons between similarly sized models, such as Google’s Gemma and Meta’s Llama, indicate distinct strengths in specific metrics, suggesting that architectural design and training methodologies also play a crucial role.

### v3

@@ Abstract
CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation
Kaiwen Yan1 *, Hongcheng Guo1 * †, Xuanqing Shi2 Shaosheng Cao3, Donglin Di2, Zhoujun Li1 1Beihang University, 2Tsinghua University, 3Xiaohongshu lin_rany@foxmail.com, hongchengguo@buaa.edu.cn, sxq23@mails.tsinghua.edu.cn caoshaosheng@xiaohongshu.com, donglin.ddl@gmail.com, lizj@buaa.edu.cn
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation. While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge Liu et al. (2024); Zhang et al. (2023). To drive progress in this field, a comprehensive and systematic evaluation framework is essential Jiang et al. (2024); Zhou et al. (2023).
This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation. Built upon insights from existing evaluation sets like McEval Chai et al. (2024) and FullStackBench Liu et al. (2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++. It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
To ensure rigorous assessment, we propose four novel evaluation metrics: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics measure models’ ability to handle multi-constraint problems from different perspectives, including full compliance, average constraint satisfaction, logical coherence, and consistency in instruction execution. By offering a structured evaluation framework, CodeIF provides valuable insights into the current state and future direction of LLM-driven code generation. Overall, our contributions are mainly four-fold:
@@
    Innovative Benchmark. We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation. CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
@@
    Automated High-Quality Instruction Generation. Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists. This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
@@
    Novel Evaluation Metrics. We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks. These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
@@
    Extensive Evaluation and Analysis. We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels. Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
@@ 2 CodeIF
As shown in Figure 1, CodeIF is constructed by collecting and refining constraint instructions from real code generation tasks, then combining these tasks with LLM outputs and human review to create a high-quality evaluation dataset.
@@ 2.1 Building
The construction of the CodeIF dataset involves two phases: collecting constraint instructions (Section 2.2) and processing data to create the final CodeIF evaluation dataset (Section 2.3).
@@ 2.2 Constraint Instructions Collection
The first phase of our work centers on code generation, constructing the CodeIF evaluation dataset through two steps: (1) collecting and verifying constraint instructions, and (2) using them for dataset generation.
We analyze benchmarks like McEval Chai et al. (2024) and FullStackBench Liu et al. (2024) to develop an instruction system spanning eight categories, each targeting specific aspects of code generation for a fine-grained assessment of LLMs’ instruction-following abilities. Constraints are decomposed into atomic instructions with explicit directives, enabling objective binary evaluation (yes/no) and minimizing subjective interpretation. The eight categories cover both architectural-level specifications and variable-level implementation controls, ensuring comprehensive constraint coverage. Specifically, the Global category evaluates adherence to overarching specifications, while Structural Control focuses on control structures (e.g., loops, conditionals) and data structures. Variable constraints assess naming and initialization. Higher abstraction levels include Interface, Function, and Class constraints for program components, while the File category tests cross-file dependencies and external library handling. The Combination category integrates constraints across dimensions, challenging models with complex scenarios.
@@
Figure 2 presents CodeIF’s distribution across programming languages and categories. The evaluation system features 8 categories and 50 fine-grained constraint instructions, systematically assessing LLMs’ code generation performance. By organizing constraints clearly, the system identifies strengths and weaknesses, guides optimization, and advances automated code generation. The full list of constraints is in Appendix 4.
@@ 2.3 Data Construction
To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval Chai et al. (2024) and FullStackBench Liu et al. (2024). These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
@@
We used large language models (LLMs) like GPT-4 to create task-specific instruction lists for code generation tasks. We prepared 20 detailed examples and formulated concise atomic instructions for accuracy. These examples guided LLMs in refining tasks and streamlining instructions to enhance clarity and output quality.
@@
We utilized LLMs to map dependencies between atomic constraints, improving our evaluation framework’s precision and verification accuracy. By understanding the dependencies among instructions, we outlined clear steps for tasks like function creation, which involve naming the function, defining parameter types, and coding the body. Incorporating these dependencies enhances our evaluation system, more accurately assessing the model’s capability with complex instructions and identifying areas for improvement. Figure 3 illustrates a CodeIF task with its instruction sequence and dependencies.
@@ 2.4 Data Analysis
Table 1 categorizes the dataset into three difficulty levels: Easy, Hard, and Full. Both Easy and Hard sets contain 600 tasks, while the Full dataset combines them, totaling 1,200 tasks across Go, Python, Java, and C++. Java has the most tasks (353), followed by Python (348), C++ (269), and Go (230). The Easy set averages 11.99 instructions per task, the Hard set 13.8, and the Full dataset 12.9, reflecting increasing complexity. Figure 4 shows task length distribution.
@@
Constraint Instruction Analysis
Table 2compares instruction distribution across difficulty levels. The Hard set consistently has more instructions per category than the Easy set, with the Global category averaging 2.5 instructions in Easy and over 3 in Hard. This indicates greater challenges for models as task complexity rises. More analysis is in Appendix D.
@@ 3 Metrics
Completely Satisfaction Rate (CSR):
@@
where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied.
Soft Satisfaction Rate (SSR):
@@
where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on.
@@
$\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$.
@@ 4.1 Experimental Setup
The temperature coefficient is set to 0 to ensure output determinism, with a maximum generation length of 4096 tokens. All other settings follow the official default parameters for each model. Commercial API models are accessed through the latest available interface as of December 2024. All experiments are conducted using the official API and 8 H800(80G).
@@ 4.2 Automatic Evaluation
To ensure robust evaluation, we used LLMs and human experts to verify model adherence to atomic constraints. Constraints were decomposed into atomic elements, enabling objective binary evaluations (Yes/No) over subjective judgments. Following FairEval Wang et al. (2023a), GPT-4-1106-Preview was the primary evaluation tool (prompt details in Appendix A). Three domain experts manually annotated 150 stratified samples. Statistical analysis showed strong agreement, with Pearson correlations of 0.87 (LLM-human) and 0.85 (inter-human), confirming high consistency. Baselines are in Appendix C.
@@ 4.3 Main Experiments
Table 3 evaluates CodeIF using four metrics: CSR, SSR, RSR, and CCSR. Detailed results are in Appendix B.
Overview.
DeepSeek-V3 and Claude-3-5-Sonnet-20241022 lead across metrics, excelling in complex tasks. However, the highest CSR on Hard tasks is just 0.362, showing challenges in meeting strict constraints.
Model Scale Trends.
Larger models generally perform better, as seen in Qwen2.5 series. However, the Llama3 series shows inconsistent results, highlighting the importance of architecture, data quality, and optimization.
Open vs. Closed Models.
Closed-source models like GPT-4O and Claude-3-5 outperform open-source models, especially under complex constraints. While large open-source models (e.g., Qwen2.5-72B-Instruct) are competitive, they lag due to differences in data quality and RLHF techniques.
Task Difficulty Impact.
Performance drops with increasing task complexity. For instance, GPT-4O’s CSR falls from 0.441 on Easy tasks to 0.325 on Hard tasks, highlighting the challenge of strict constraints.
@@ 5.1 Performance Analysis Across Instruction Types
Figure 5 compares LLM performance across instruction categories, revealing notable variations. DeepSeek-V3 leads overall, excelling in combination tasks (0.831) and global structure control, though weaker in variable handling, reflecting its optimization focus. Meta’s Llama series shows a clear correlation between model size and performance, with larger variants (Llama-3.3-70B-Instruct) outperforming smaller ones (Llama-3.2-1B-Instruct). However, size alone is not decisive; comparisons with similarly sized models like Google’s Gemma highlight the role of architecture and training methods in shaping performance.